Two point charges +4e and +e are fixed at a
distance 'a'. A third charge q is placed on a
straight line joining these two charges so that
q is in equilibrium. Determine the position
of q.
1
Expert's answer
2019-07-04T11:10:43-0400

Let q be r distant from e



Force on q due to e 


"=k\frac{qe}{r^2}"

 

Force on q due to 4e 


"=k\frac{q4e}{(a-r)^2}"


For it to be in equilibrium : 

The forces must be equal and opposite. 


"k\frac{qe}{r^2}=k\frac{4eq}{(a-r)^2}"

"\implies \frac{1}{r^2}=\frac{4}{(a-r)^2}"


"\implies {a^2}+{r^2}-2{a}{r}=4{r^2}"


"\implies 3{r^2}+2{a}{r}-{a^2}=0"

 


"\implies 3{r^2}+3{a}{r}-{a}{r}-{a^2}=0"


"\implies 3r(r+a)-a(r+a)=0"

 

"\implies (3r-a)(r+a)=0"

"\implies r=-a" or "r = \frac{a}{3}"

Since distance is non-negative 

"r = \frac{a}{3}" should be the answer .